A nickel wire has a resistance of 20 ohms at 10 degrees Celsius. After several hours of heating, the resistance of the coil has increased to 26.38 ohms. What is the new temperature of the wire?


temperature coefficient of nickel - 0.006


1
Expert's answer
2021-02-22T07:15:10-0500

We have given,


"R_o=20\Omega\\

T_o=10^{\circ} C\\

R_t=26.38\Omega"


Let T be the new temperature of the wire

We know,


"R_t=R_o(1+\alpha\Delta T)"


Putting the values of "R_o,T_o and R_t"


"26.38=20(1+0.006(T_o-10))"


"53.166=T-10"


"T=63.16^\circ C"


Hence,the new temperature of wire is"63.16^\circ C"